# Rustam Ibragimova,b    Jihyun Kimc    Anton Skrobotovd a Imperial College Business School b New Economic School

### v2

@@ 1 Introduction
Predictive regressions play a central role in empirical finance, providing a framework for assessing whether financial or macroeconomic variables can forecast future returns. Prominent applications include the forecasting of equity and aggregate returns (see, among others, CampbellYogo2006, CampbellYogo2006; GoyalWelch, GoyalWelch; campbell, campbell, Hirshleifer, Hirshleifer; KJ, KJ; Rapach, Rapach; MR, MR; Goyal2024, Goyal2024, and references therein) and tests of market efficiency (e.g., Fama1, Fama1, Fama, Fama2, the review in Martin, Martin, and references therein). Numerous studies have examined the econometric properties of predictive regressions for stock returns (see Phillips2015, Phillips2015, for a review), highlighting several statistical challenges that complicate inference on return predictability. Widely used predictors, including valuation ratios such as the dividend–price and earnings–price ratios, typically exhibit near–unit–root dynamics, and their innovations are correlated with stock returns over long horizons. This combination of persistence and endogeneity induces substantial biases in conventional hypothesis tests (see, e.g., stambaugh1999predictive, stambaugh1999predictive; KimPark1, KimPark1). Moreover, stock return volatility is stochastic and highly persistent (jacquier2004bayesian, jacquier2004bayesian; hansen2014estimating, hansen2014estimating), and cavaliere2004testing shows that such volatility can lead to severe size distortions in tests assuming stationarity. Predictive regression data also often exhibit heavy tails, jumps, structural breaks, and regime shifts, further undermining standard inference (op. cit.).
@@
A large literature has addressed persistent endogeneity in predictive regressions. Notably, CampbellYogo2006, ChenDeo2009, PhillipsMagdalinos2009, and KMS2015, among others, proposed inference procedures designed to handle persistence and endogeneity. While these methods perform well in such environments, they do not adequately address other empirical features such as persistent volatility, structural breaks, or regime switching. The simulations in ibragimov2023new show that standard tests can suffer from severe size distortions under persistent volatility.
@@
CJP2016 proposed an inference method (the Cauchy RT) based on the Cauchy estimator and a time-change transformation in a continuous-time framework to address endogenous persistence and volatility in the errors. [^2] See also BuKimWang for an alternative method robust to endogenously persistent or heavy-tailed regressors and persistent volatility in continuous time. ibragimov2023new introduced another approach (the Cauchy VC), also based on the Cauchy estimator but with a nonparametric volatility correction. Unlike CJP2016, it applies to both continuous- and discrete-time models.
@@
This paper proposes two practical tests that serve as robust alternatives to these methods. The proposed tests are robust to heterogeneous and persistent volatility, as well as to endogenous, persistent, and/or heavy-tailed regressors. Both employ Cauchy estimation procedures, as in CJP2016 and ibragimov2023new, to address endogeneity, persistence, and heavy tails. The two methods differ in their treatment of heterogeneous volatility: the first extends the $t$-statistic-based group inference of IbragimovMuller2010 to asymptotically normal Cauchy estimators, while the second is a hybrid test combining Cauchy and OLS estimation that uses the Cauchy estimator for the coefficient and OLS residuals for the standard error.
@@
The proposed methods are easy to implement and avoid the technical complexities of the time-change transformation in CJP2016 and the nonparametric volatility correction in ibragimov2023new. Although they rely on an asymptotically exogenous volatility condition, they perform well in finite samples and compare favorably to alternative procedures across empirically relevant scenarios, including mild violations of this condition. Moreover, the proposed methods apply to both continuous- and discrete-time models, as in ibragimov2023new. Overall, the two tests developed here and the inference procedures of CJP2016 and ibragimov2023new should be viewed as complementary, particularly in environments with heteroskedastic and persistent volatility.
@@
The remainder of the paper is organized as follows. Section 2 discusses the predictive regression model and the Cauchy estimator; Section 3 develops the inference procedures and theoretical results; Section 4 extends them to multiple predictors and models with intercepts; Sections 5 and 6 present simulations and an empirical illustration; and Section 7 concludes. All proofs are in the Appendix.
@@ 2.1 Model and Issues
Conditions (a) and (b) are standard and ensure that the conditional variance of $u_{t}$ is identified: $E(u_{t}^{2}|\mathcal{F}_{t-1})=v_{t}^{2}$. Condition (c) is a conditional Lindeberg condition, which holds, for example, if $\sup_{1\leq t\leq T}E(|\varepsilon_{t}|^{2+\delta}|\mathcal{F}_{t-1})$ is bounded for some $\delta>0$ with probability one. See ibragimov2023new and references therein for further discussion and examples of processes satisfying Assumption 2.1.
The hypothesis of unpredictability of $(y_{t})$ corresponds to $H_{0}:\beta=0$ in regression (1). It is well-known that standard OLS $t$-statistic inference is not robust to many empirically relevant features of financial data. For instance, the OLS estimator of $\beta$ is not asymptotically Gaussian under $H_{0}$ if $(x_{t})$ is endogenous and (nearly) nonstationary (see elliott1994inference, elliott1994inference; PhillipsNear, PhillipsNear; GP, GP; PM, PM; KMS2015, KMS2015), or even if $(x_{t})$ is stationary but has infinite variance (e.g., GrangerOrr, GrangerOrr; EKM, EKM; ibragimov2015heavy, ibragimov2015heavy). This non-Gaussianity persists even when the errors are homoskedastic with $v_{t}^{2}=\sigma^{2}$ for all tt. [^3] As usual, the endogeneity of $(x_{t-1})$ refers to the existence of nonzero long-run covariance between the innovations of $(y_{t})$ and $(x_{t-1})$. Furthermore, stock return data exhibit time-varying and stochastically persistent volatility, which causes the distribution of the OLS $t$-statistic to deviate from standard normality, leading to size distortions in conventional tests (see CJP2016, CJP2016; ibragimov2023new, ibragimov2023new).
@@ 2.2 The Cauchy Estimator
where $\text{sign}(\cdot)$ denotes the sign function, $\text{sign}(x)=1$ for $x\geq 0$ and $\text{sign}(x)=-1$ for $x<0$. The estimator $\check{\beta}$ can be interpreted as an instrumental variable (IV) estimator using $\text{sign}(x_{t-1})$ as an instrument for $x_{t-1}$ (see, e.g., SoShin, SoShin; BREITUNG2015, BREITUNG2015; kim-meddahi-2019, kim-meddahi-2019; shephard2020, shephard2020).
@@
which takes values in $\mathbf{D}_{\mathbb{R}}[0,1]$, the space of càdlàg functions on $[0,1]$ with values in $\mathbb{R}^{d}$. By the functional central limit theorem for martingales (Theorem 18.2 of billingsley1986convergence, billingsley1986convergence), we have $W^{T}\Rightarrow W$ in $\mathbf{D}_{\mathbb{R}}[0,1]$, where $W$ is a standard Brownian motion.
@@
Following ibragimov2023new, we assume that the volatility process $\sigma^{T}$ is persistent in the sense that it admits a limiting process $\sigma$ defined on $[0,1]$ such that $(W^{T},\sigma^{T})\Rightarrow(W,\sigma)$ jointly.
@@
Assumption 2.2 encompasses a wide class of models, including those with nonstationary volatility, regime switching, or structural breaks. [^4] Assumptions 2.1 and 2.2 exclude some globally homoskedastic processes, such as stationary GARCH models. However, the hybrid testing procedure proposed later remains valid under $T^{-1}\sum_{t=1}^{T}v_{t}^{2}\to_{p}\omega^{2}>0$, which includes conditionally heteroskedastic but globally homoskedastic processes, such as stationary GARCH models (see also Section 4 of ibragimov2023new, ibragimov2023new). It also covers cases with deterministic volatility $v_{t}=\sigma(t/T)$, as in cavaliere-taylor-2007, cavaliere-taylor-2008, xu-phillips-2008, and harvey-leybourne-zu, among others. [^5] Assumption 2.2 is a simplified version of the condition $v_{[Tr]}/a_{T}\Rightarrow\sigma_{r}$ in Assumption 2 of cavaliere-taylor-2009. We focus on stochastically bounded volatilities with $a_{T}=1$, excluding explosive volatility settings ($a_{T}\to\infty$) for simplicity. It further includes nonstationary volatility processes such as those in hansen1995regression and chung2007nonstationary, where $v_{t}^{2}$ is a smooth positive transformation of a (near) unit root process. Overall, Assumptions 2.1 and 2.2 are general enough to allow for stochastic and discontinuous volatility—features commonly observed in financial returns.
@@
by standard results on the convergence of stochastic integrals (see hansen-1992, hansen-1992; kurtz1991weak, kurtz1991weak; ibragimov2023new, ibragimov2023new). The limit $\int_{0}^{1}\sigma(r)\,dW(r)$ is in general a non-Gaussian martingale, becoming Gaussian only if $W$ and $\sigma$ are independent. In that case, $\int_{0}^{1}\sigma(r)\,dW(r)$ is a scale mixture of normals with variance $\int_{0}^{1}\sigma^{2}(r)\,dr$, denoted
@@ 3.1 Robust $t$-Statistic Approach
The first approach relies on $t$-statistic-based inference using group estimates of $\beta$, as proposed by IbragimovMuller2010 (see also IM1, IM1; Section 3.3 of ibragimov2015heavy, ibragimov2015heavy). The method is based on normalized Cauchy estimators—specifically, the numerator of the Cauchy estimator divided by $\sqrt{T}$ in the full-sample case:
@@
The statistics $\{\check{\gamma}_{j}\}_{j=1}^{q}$ do not satisfy the standard condition in IbragimovMuller2010, which requires estimators $\{\tilde{\beta}_{j}\}_{j=1}^{q}$ such that
@@
Proposition 3.2 shows that the $t$-statistic approach is conservative under $H_{0}$ and consistent under $H_{A}$ when $(x_{t})$ is stationary with a finite first moment. It is thus valid and robust to persistent heteroskedasticity and endogenously heavy-tailed covariates. However, it becomes inconsistent for highly persistent covariates, such as (near) unit-root processes. To illustrate, consider the generalized local-to-unity framework of dou2021generalized, where $X^{T}(r)=x_{[Tr]}$ for $r\in[0,1]$ and
@@
with $X$ a stationary continuous-time Gaussian ARMA process. [^6] See dou2021generalized for a detailed discussion.
@@ 3.2 A Hybrid Test
A key feature of the Cauchy estimator $\check{\beta}$ is that its properly normalized limit distribution is invariant to the data-generating process of $(x_{t})$. By contrast, the OLS estimator’s variance depends on both $(u_{t})$ and $(x_{t})$, complicating variance estimation even under homoskedasticity. For $\check{\beta}$, the asymptotic variance depends solely on $u_{t}$, requiring only heteroskedasticity-robust adjustments. [^8] See shephard2020, Section 4.3, for related discussion.
@@
Here, $\hat{\omega}^{2}$ estimates $\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$ using OLS residuals. As noted by shephard2020, the Cauchy-based variance estimator performs poorly because the Cauchy estimator converges more slowly and less efficiently than OLS when $(x_{t})$ is heavy-tailed or nearly integrated. Hence, we use OLS residuals to improve efficiency. [^9] A related approach is employed by KMS2015 in the IVX framework of PM.
@@
Assumption 3.1 is very general and holds in many time-series settings. It is weaker than Assumption 3.2 of ibragimov2023new, which requires $\sum_{t=1}^{T}x_{t-1}u_{t}=O_{p}\!\left(\sqrt{T^{p}\sum_{t=1}^{T}x_{t-1}^{2}}\right)$ for $p\in[0,1/16)$. As shown in ibragimov2023new, this holds with $p=0$ when $(x_{t})$ is either (near) unit root or stationary with finite variance; it also applies to certain stationary heavy-tailed processes (see, e.g., Samorodnitsky2007, Samorodnitsky2007).
@@
Remark. Proposition 3.4(a) also holds if $\tau(\check{\beta})$ uses $\bar{\omega}^{2}=T^{-1}\sum_{t=1}^{T}y_{t}^{2}$ instead of $\hat{\omega}^{2}$, since $\beta=0$ under $H_{0}$. Moreover, the corresponding test remains consistent when $(x_{t})$ is stationary with finite variance or follows a generalized local-to-unity process (dou2021generalized). However, it can be inconsistent for heavy-tailed $(x_{t})$. For instance, if $(x_{t})$ is i.i.d. $\alpha$-stable with $\alpha\in(0,2)$ and independent of $(u_{t})$, then
@@
by the generalized central limit theorem (see feller1971, feller1971; logan1973, logan1973; Davis1983, Davis1983; DavisResnick1986, DavisResnick1986). Thus, the use of $\hat{\omega}^{2}$ (or another consistent estimator under both $H_{0}$ and $H_{A}$) is crucial for the consistency of the hybrid test.

### v3

# Rustam Ibragimova,b    Jihyun Kimd    Anton Skrobotove a Imperial College Business School b New Economic School

@@ 1 Introduction
Predictive regressions play a central role in empirical finance, providing a framework for assessing whether financial or macroeconomic variables can forecast future returns. Prominent applications include the forecasting of equity and aggregate returns (see, among others, Campbell and Yogo, 2006; Goyal and Welch, 2006; Campbell and Thompson, 2008; Hirshleifer et al., 2009; Kelly and Jiang, 2014; Rapach et al., 2016; Møller and Rangvid, 2015; Goyal et al., 2024, and references therein) and tests of market efficiency (e.g., Fama, 1970, 1984, 1991; the review in Martin and Nagel, 2022, and references therein). A large body of work has examined the econometric properties of predictive regressions (see Phillips, 2015 for a review), highlighting several statistical challenges that complicate inference on return predictability.
@@
A key difficulty arises from the persistence and endogeneity of commonly used predictors. Variables such as the dividend–price and earnings–price ratios typically exhibit near–unit–root behavior, and their innovations are correlated with future returns over long horizons. This combination induces substantial biases in conventional hypothesis tests (see, e.g., Stambaugh, 1999; Kim and Park, 2017). In addition, stock return volatility is stochastic and highly persistent (Jacquier et al., 2004; Hansen and Lunde, 2014), and Cavaliere (2004) shows that such volatility can lead to severe size distortions in tests that rely on stationarity. Predictive regression data also often exhibit heavy tails, jumps, structural breaks, and regime switching, further undermining standard inference methods.
@@
A large literature has proposed procedures to address persistent endogeneity in predictive regressions. Notably, Campbell and Yogo (2006), Chen and Deo (2009), Phillips and Magdalinos (2009), and Kostakis et al. (2015), among others, develop inference methods that perform well in the presence of persistent and endogenous regressors. However, these approaches typically rely on assumptions such as homoskedasticity or weak dependence in volatility and may not remain valid under empirically relevant features such as persistent or nonstationary volatility, structural breaks, or regime switching. For example, Ibragimov et al. (2024) show that standard tests can suffer from severe size distortions in the presence of persistent volatility.
@@
To address these limitations, Choi et al. (2016) propose an inference method (the Cauchy RT) based on the Cauchy estimator and a time-change transformation in a continuous-time framework. [^2] See also Bu et al. (2023) for an alternative method robust to endogenously persistent or heavy-tailed regressors and persistent volatility in continuous time. Ibragimov et al. (2024) introduce another approach (the Cauchy VC), also based on the Cauchy estimator but with a nonparametric volatility correction, which applies to both continuous- and discrete-time models.
@@
This paper proposes two practical tests that serve as robust alternatives to these methods. The proposed procedures are robust to heterogeneous and persistent volatility, as well as to endogenous, persistent, and heavy-tailed regressors. Both methods employ Cauchy estimation, as in Choi et al. (2016) and Ibragimov et al. (2024), to address endogeneity, persistence, and heavy tails. The two approaches differ in their treatment of heterogeneous volatility: the first extends the $t$-statistic–based group inference of Ibragimov and Müller (2010) to asymptotically normal Cauchy estimators, while the second is a hybrid procedure that combines Cauchy and OLS estimation by using the Cauchy estimator for the coefficient and OLS residuals for variance estimation.
@@
The proposed methods are simple to implement and avoid the technical complexities associated with existing approaches, such as the time-change transformation in Choi et al. (2016) or the nonparametric volatility correction in Ibragimov et al. (2024). Although they rely on an asymptotic exogeneity condition on volatility, simulation results show that they perform well in finite samples, even under mild violations of this condition. Moreover, the proposed methods apply to both continuous- and discrete-time models. Overall, our procedures complement existing Cauchy-based methods and provide a practical alternative for robust inference under a broad range of empirically relevant environments.
In addition to the continuous-time framework, we evaluate the performance of our methods in discrete-time predictive regressions and compare them with the IVX approach of Kostakis et al. (2015). The IVX method is specifically designed to address persistence and endogeneity and performs well under homoskedastic or weakly dependent volatility. Our results highlight that the two approaches are complementary. While IVX delivers strong performance under standard conditions, it may suffer from size distortions in the presence of nonstationary or persistent volatility. By contrast, our methods are designed to remain robust to such features, including heavy tails and time-varying volatility. This comparison underscores the advantages of each approach across empirically relevant settings.
@@
The remainder of the paper is organized as follows. Section 2 introduces the predictive regression model and the Cauchy estimator. Section 3 develops the proposed inference procedures and establishes their theoretical properties. Section 4 extends the analysis to models with multiple predictors and intercepts. Sections 5 and 6 present simulation results and an empirical application. Section 7 concludes. All proofs are provided in the Appendix, and additional theoretical analysis, as well as supplementary simulation and empirical results, are reported in the accompanying Online Appendix.
@@ 2.1 Model and Issues
Conditions (a) and (b) are standard and ensure that the conditional variance of $u_{t}$ is identified: $E(u_{t}^{2}|\mathcal{F}_{t-1})=v_{t}^{2}$. Condition (c) is a conditional Lindeberg condition, which holds, for example, if $\sup_{1\leq t\leq T}E(|\varepsilon_{t}|^{2+\delta}|\mathcal{F}_{t-1})$ is bounded for some $\delta>0$ with probability one. See Ibragimov et al. (2024) and references therein for further discussion and examples of processes satisfying Assumption 2.1.
The hypothesis of unpredictability of $(y_{t})$ corresponds to $H_{0}:\beta=0$ in regression (1). It is well-known that standard OLS $t$-statistic inference is not robust to many empirically relevant features of financial data. For instance, the OLS estimator of $\beta$ is not asymptotically Gaussian under $H_{0}$ if $(x_{t})$ is endogenous and (nearly) nonstationary (see Elliott and Stock, 1994; Phillips, 1987; Giraitis and Phillips, 2006; Phillips and Magdalinos, 2007a; Kostakis et al., 2015), or even if $(x_{t})$ is stationary but has infinite variance (e.g., Granger and Orr, 1972; Embrechts et al., 1997; Ibragimov et al., 2015). This non-Gaussianity persists even when the errors are homoskedastic with $v_{t}^{2}=\sigma^{2}$ for all tt. [^3] As usual, the endogeneity of $(x_{t-1})$ refers to the existence of nonzero long-run covariance between the innovations of $(y_{t})$ and $(x_{t-1})$. Furthermore, stock return data exhibit time-varying and stochastically persistent volatility, which causes the distribution of the OLS $t$-statistic to deviate from standard normality, leading to size distortions in conventional tests (see Choi et al., 2016; Ibragimov et al., 2024).
@@ 2.2 The Cauchy Estimator
where $\text{sign}(\cdot)$ denotes the sign function, $\text{sign}(x)=1$ for $x\geq 0$ and $\text{sign}(x)=-1$ for $x<0$. The estimator $\check{\beta}$ can be interpreted as an instrumental variable (IV) estimator using $\text{sign}(x_{t-1})$ as an instrument for $x_{t-1}$ (see, e.g., So and Shin, 1999; Breitung and Demetrescu, 2015; Kim and Meddahi, 2020; Shephard, 2020).
@@
which takes values in $\mathbf{D}_{\mathbb{R}}[0,1]$, the space of càdlàg functions on $[0,1]$ with values in $\mathbb{R}^{d}$. By the functional central limit theorem for martingales (Theorem 18.2 of Billingsley, 1986), we have $W^{T}\Rightarrow W$ in $\mathbf{D}_{\mathbb{R}}[0,1]$, where $W$ is a standard Brownian motion.
@@
Following Ibragimov et al. (2024), we assume that the volatility process $\sigma^{T}$ is persistent in the sense that it admits a limiting process $\sigma$ defined on $[0,1]$ such that $(W^{T},\sigma^{T})\Rightarrow(W,\sigma)$ jointly.
@@
Assumption 2.2 encompasses a wide class of models, including those with nonstationary volatility, regime switching, or structural breaks. [^4] Assumptions 2.1 and 2.2 exclude some globally homoskedastic processes, such as stationary GARCH models. However, the hybrid testing procedure proposed later remains valid under $T^{-1}\sum_{t=1}^{T}v_{t}^{2}\to_{p}\omega^{2}>0$, which includes conditionally heteroskedastic but globally homoskedastic processes, such as stationary GARCH models (see also Section 4 of Ibragimov et al., 2024). It also covers cases with deterministic volatility $v_{t}=\sigma(t/T)$, as in Cavaliere and Taylor (2007, 2008), Xu and Phillips (2008), and Harvey et al. (2019), among others. [^5] Assumption 2.2 is a simplified version of the condition $v_{[Tr]}/a_{T}\Rightarrow\sigma_{r}$ in Assumption 2 of Cavaliere and Taylor (2009). We focus on stochastically bounded volatilities with $a_{T}=1$, excluding explosive volatility settings ($a_{T}\to\infty$) for simplicity. It further includes nonstationary volatility processes such as those in Hansen (1995) and Chung and Park (2007), where $v_{t}^{2}$ is a smooth positive transformation of a (near) unit root process. Overall, Assumptions 2.1 and 2.2 are general enough to allow for stochastic and discontinuous volatility—features commonly observed in financial returns.
@@
by standard results on the convergence of stochastic integrals (see Hansen, 1992; Kurtz and Protter, 1991; Ibragimov et al., 2024). The limit $\int_{0}^{1}\sigma(r)\,dW(r)$ is in general a non-Gaussian martingale, becoming Gaussian only if $W$ and $\sigma$ are independent. In that case, $\int_{0}^{1}\sigma(r)\,dW(r)$ is a scale mixture of normals with variance $\int_{0}^{1}\sigma^{2}(r)\,dr$, denoted
@@ 3.1 Robust $t$-Statistic Approach
The first approach relies on $t$-statistic-based inference using group estimates of $\beta$, as proposed by Ibragimov and Müller (2010) (see also Ibragimov and Müller, 2016; Section 3.3 of Ibragimov et al., 2015). The method is based on normalized Cauchy estimators—specifically, the numerator of the Cauchy estimator divided by $\sqrt{T}$ in the full-sample case:
@@
The statistics $\{\check{\gamma}_{j}\}_{j=1}^{q}$ do not satisfy the standard condition in Ibragimov and Müller (2010), which requires estimators $\{\tilde{\beta}_{j}\}_{j=1}^{q}$ such that
@@
Proposition 3.2 shows that the $t$-statistic approach is conservative under $H_{0}$ and consistent under $H_{A}$ when $(x_{t})$ is stationary with a finite first moment. It is thus valid and robust to persistent heteroskedasticity and endogenously heavy-tailed covariates. However, it becomes inconsistent for highly persistent covariates, such as (near) unit-root processes. To illustrate, consider the generalized local-to-unity framework of Dou and Müller (2021), where $X^{T}(r)=x_{[Tr]}$ for $r\in[0,1]$ and
@@
with $X$ a stationary continuous-time Gaussian ARMA process. [^6] See Dou and Müller (2021) for a detailed discussion.
@@ 3.2 A Hybrid Test
A key feature of the Cauchy estimator $\check{\beta}$ is that its properly normalized limit distribution is invariant to the data-generating process of $(x_{t})$. By contrast, the OLS estimator’s variance depends on both $(u_{t})$ and $(x_{t})$, complicating variance estimation even under homoskedasticity. For $\check{\beta}$, the asymptotic variance depends solely on $u_{t}$, requiring only heteroskedasticity-robust adjustments. [^8] See Shephard (2020), Section 4.3, for related discussion.
@@
Here, $\hat{\omega}^{2}$ estimates $\omega^{2}=\int_{0}^{1}\sigma^{2}(r)\,dr$ using OLS residuals. As noted by Shephard (2020), the Cauchy-based variance estimator performs poorly because the Cauchy estimator converges more slowly and less efficiently than OLS when $(x_{t})$ is heavy-tailed or nearly integrated. Hence, we use OLS residuals to improve efficiency. [^9] A related approach is employed by Kostakis et al. (2015) in the IVX framework of Phillips and Magdalinos (2007a).
@@
Assumption 3.1 is very general and holds in many time-series settings. It is weaker than Assumption 3.2 of Ibragimov et al. (2024), which requires $\sum_{t=1}^{T}x_{t-1}u_{t}=O_{p}\!\left(\sqrt{T^{p}\sum_{t=1}^{T}x_{t-1}^{2}}\right)$ for $p\in[0,1/16)$. As shown in Ibragimov et al. (2024), this holds with $p=0$ when $(x_{t})$ is either (near) unit root or stationary with finite variance; it also applies to certain stationary heavy-tailed processes (see, e.g., Samorodnitsky et al., 2007).
@@
Remark. Proposition 3.4(a) also holds if $\tau(\check{\beta})$ uses $\bar{\omega}^{2}=T^{-1}\sum_{t=1}^{T}y_{t}^{2}$ instead of $\hat{\omega}^{2}$, since $\beta=0$ under $H_{0}$. Moreover, the corresponding test remains consistent when $(x_{t})$ is stationary with finite variance or follows a generalized local-to-unity process (Dou and Müller, 2021). However, it can be inconsistent for heavy-tailed $(x_{t})$. For instance, if $(x_{t})$ is i.i.d. $\alpha$-stable with $\alpha\in(0,2)$ and independent of $(u_{t})$, then
@@
by the generalized central limit theorem (see Feller, 1971; Logan et al., 1973; Davis, 1983; Davis and Resnick, 1986). Thus, the use of $\hat{\omega}^{2}$ (or another consistent estimator under both $H_{0}$ and $H_{A}$) is crucial for the consistency of the hybrid test.